# Detection and Reconstruction of a Random Hypergraph from Noisy Graph Projection

### v1

@@ Abstract
For a $d$-uniform random hypergraph on $n$ vertices in which hyperedges are included i.i.d. so that the average degree in the hypergraph is $n^{\delta+o(1)}$, the projection of such a hypergraph is a graph on the same $n$ vertices where an edge connects two vertices if and only if they belong to a same hyperedge. In this work, we study the inference problem where the observation is a noisy version of the graph projection where each edge in the projection is kept with probability $p=n^{-1+\alpha+o(1)}$ and each edge not in the projection is added with probability $q=n^{-1+\beta+o(1)}$. For all constant $d$, we establish sharp thresholds for both detection (distinguishing the noisy projection from an Erdős-Rényi random graph with edge density $q$) and reconstruction (estimating the original hypergraph). Notably, our results reveal a detection-reconstruction gap phenomenon in this problem. Our work also answers a problem raised in [BGPY25].
@@ 1 Introduction
In this paper, we address the problem of reconstructing the multi-vertex associations of a hypergraph from its projected graph, a graph in which two vertices are connected if and only if they belong to the same hyperedge. More precisely, we pose the reconstruction task as follows: for any constant $d\in\mathbb{N}$, we denote by $\operatorname{U}=\operatorname{U}_{n,d}$ the collection containing all subsets of $[n]=\{1,\ldots,n\}$ with cardinality $d$.
@@
Define $\mathcal{H}$ to be the random $d$-uniform hypergraph on $[n]$ generated by including each $\Psi\in\operatorname{U}_{n,d}$ in $\mathcal{H}$ with probability $s=n^{-(d-1)+\delta+o(1)}$ for some constant $10\leq\delta<$ $1$For any two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write equivalently $a_{n}=O(b_{n})$, $b_{n}=\Omega(a_{n})$, $a_{n}\lesssim b_{n}$ and $b_{n}\gtrsim a_{n}$ if there exists a positive absolute constant $c$ such that $a_{n}/b_{n}\leq c$ holds for all $n$. We write $a_{n}=o(b_{n})$, $b_{n}=\omega(a_{n})$, $a_{n}\ll b_{n}$, and $b_{n}\gg a_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. We write $a_{n}=\Theta(b_{n})$ if both $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$ hold. Without further specification, all asymptotics are taken with respect to $n\to\infty$.. We denote the law of $\mathcal{H}$ by $\mu=\mu_{n,d,\delta}$. For simplicity, we will also use $\mathcal{H}$ to denote the set of hyperedges and $|\mathcal{H}|$ the cardinality of $\mathcal{H}$. In addition, for any $d$-uniform hypergraph $\mathcal{H}$, define $\mathcal{P}(\mathcal{H})$ to be the projection graph on $[n]$ where for all unordered pairs $(i,j)$ with $i\neq j$
@@
We require the condition $q\leq p$, reflecting the natural expectation that elements of $\mathcal{P}(\mathcal{H})$ tend to appear in the graph $A$. Inspired by real-world social networks such as scientific co-authorship networks [New04], graphs with group structure (e.g., random intersection graph [BBN20, vdHKV21] and hypergraph stochastic block model [GD17, ACKZ15, SZ22, GJ23, DWar, GP23, GP24]) have garnered significant attention in the research community. Among them, the model of noisy random hypergraph projection serves as a natural toy model to study how to infer the original group structure given only pairwise vertex information. For a more comprehensive overview on this model and its various applications in real networks, we refer the readers to [BGPY25]. Two natural inference problems regarding this model are as follows: (1) the detection problem, i.e., testing $\mathbb{P}$ against $\mathbb{Q}=\mathbb{Q}_{n,q}$, where $\mathbb{Q}$ is the law of an Erdős-Rényi graph on $[n]$ with edge density $q=n^{-1+\beta+o(1)}$; (2) the reconstruction problem, i.e., estimating the original hypergraph $\mathcal{H}$ according to the observation $A(\mathcal{H})$.
The problem of reconstruction from the projected graph has been extensively studied in recent years, both theoretically and empirically. For instance, [WK24] assumes access to a sample of a hypergraph from the underlying distribution and proposes a scoring method to select hyperedges based on their similarity to the sampled hypergraph. Another Bayesian algorithmic approach (i.e., to sample from the posterior distribution given the projected graph) was proposed in [YPP21, LYA23]. In addition, in [CFLL24] a large foundational model is used to recover a weighted hypergraph from a sample of its hyperedges, where each hyperedge is assigned a probability proportional to its weight. In a series of recent progress [BGP24, BGPY25], the sharp threshold for reconstructing the hypergraph $\mathcal{H}$ based on $A(\mathcal{H})$ in the noiseless setting $p=1,q=0$ was determined. In particular, it was shown in [BGPY25] that when $\delta>\tfrac{d-1}{d+1}$ it is information theoretically impossible to reconstruct a positive fraction of $\mathcal{H}$ and when $\delta<\tfrac{d-1}{d+1}$ there exists an efficient algorithm that reconstructs a $1-o(1)$ fraction of $\mathcal{H}$. They further raised the question of whether an analogous reconstruction threshold can be established in the noisy setting. Our main result can be informally summarized as follows:
@@
    (Almost-exact reconstruction) Suppose either $p=1-\Omega(1)$, $\delta>\tfrac{d-1}{d+1}$ or $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then it is informationally impossible to reconstruct a $1-o(1)$ fraction of $\mathcal{H}$. Conversely, if $p=1-o(1),\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$, then there exists an efficient algorithm that successfully reconstructs a $1-o(1)$ fraction of $\mathcal{H}$.
Notably, our result reveals a detection-reconstruction gap phenomenon in this problem where when $p=\Theta(1)$ (which implies that $\alpha=1$), detection is possible for all $\beta,\delta\in(0,1)$ but (partial) reconstruction is possible if and only if $\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. See Figure 1 for a summary of the regions where detection/reconstruction is possible. The reader may wonder whether the detection-recovery gap could be eliminated by choosing an alternative null distribution whose edge density matched that of the planted model $\mathbb{P}$. We show in Appendix A that this adjustment closes the gap partially but not entirely. In particular, detection remains possible at least whenever $\delta,\beta<\tfrac{d-1}{d+1}$, noting that for $\delta<\tfrac{d-1}{d+1}$ we always have $\tfrac{d-1}{d+1}>\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$.
Figure 1: Summary of the detection-reconstruction gap when $p=\Theta(1)$. Red regime $I$: reconstruction of $\mathcal{H}$ is possible; Blue regime $II$: detection with a better null distribution $\widetilde{\mathbb{Q}}$ (that matches the average edge-density) is possible, but reconstruction of $\mathcal{H}$ is impossible; Yellow regime $III$: detection with the simple null distribution $\mathbb{Q}$ is possible, but detection with a better null distribution $\widetilde{\mathbb{Q}}$ is unknown.
@@ 1.2 The reconstruction threshold
Recall the definition of $s$ in 1.1. A test statistic ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ is said to achieve
@@
^:={${}\widehat{\mathcal{H}}:=\big{\{}\Psi\in\operatorname{U}:A_{i,j}=1\mbox{ for all }i,j\in\Psi\big{\}}\,.$ (1.4)
@@
Suppose $\delta<\tfrac{d-1}{d+1},\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$ and $p=1-o(1)$. We have
@@
The proof of Proposition 1.7 is incorporated in Section 4. Next we show the impossibility of weak reconstruction. Note that when $\delta>\tfrac{d-1}{d+1}$, using [BGPY25, Theorem 1.1] it is informationally impossible to achieve weak reconstruction even when $p=1,q=0$. Thus, without loss of generality, we assume that $\delta\leq\tfrac{d-1}{d+1}$.
@@
    Suppose that $p=o(1)$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
    Suppose $q=n^{-1+\beta+o(1)}$ where $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
Suppose that $p=1-\Omega(1)$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
[…]
@@
Let $M_{k}(\mathcal{H})$ be the set of $(i,j)\in\mathcal{P}(\mathcal{H})$ such that there exists at least $k$ hyperedges in $\mathcal{H}$ that contain $(i,j)$. In addition, define
@@
which yields (2.2). Now conditioning on $\mathcal{P}(\mathcal{H})$ satisfying (2.2), by (2.1) and a standard binomial tail estimate, we obtain that with probability $1-o(1)$ under $\mathbb{P}$
@@
Based on Lemma 3.2, it suffices to bound the tail probability of $|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|$, as incorporated in the following lemma.
@@
The proof of inequality can be found in, e.g., [AS16, Section 7.7]. We now proceed to the proof of Lemma 3.3.
@@
where we view $(\mathcal{H},\mathcal{H}^{\prime})$ as a vector in $\{0,1\}^{2\binom{n}{d}}$. It is straightforward to verify that $h$ is Lipschitz. Furthermore, we claim that the function $h$ defined above is $f$-certifiable, where $f(s)=2\tbinom{d}{2}s\leq d^{2}s$. In fact, if $h(\mathcal{H},\mathcal{H}^{\prime})\geq s$ for some $(\mathcal{H},\mathcal{H}^{\prime})\in\Omega$, then there exists $E=\{(i,j):1\leq i<j\leq n\}$ such that for all $e\in E$ there exist $\zeta_{e},\zeta_{e}^{\prime}\in\tbinom{[n]}{d}$ with $\mathcal{H}(\zeta_{e})=\mathcal{H}^{\prime}(\zeta^{\prime}_{e})=1$. Denote $I=\{\zeta_{e}:e\in E\}\cap\{\zeta^{\prime}_{e}:e\in E\}$. Then for all (~,~)$(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\in\{0,1\}^{2\binom{n}{d}}$ that agrees with $(\mathcal{H},\mathcal{H}^{\prime})$ on $I$, we have $e\in E(\mathcal{P}(\widetilde{\mathcal{H}}))\cap E(\mathcal{P}(\widetilde{\mathcal{H}}^{\prime}))$ and thus $h(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\geq s$. As a result, Theorem 3.5 yields that for any $b,t\geq 0$,
$\mathbb{P}\big{(}X\leq b-td\sqrt{2b}\big{)}\cdot\mathbb{P}(X\geq b)\leq e^{-t^{2}/4}\,.$
@@
$\mathbb{P}\Big{(}X\leq\operatorname{Median}(X)-t\Big{)}\leq 2e^{-t^{2}/4d\operatorname{Median}(X)}\,.$
In addition, taking $b=\operatorname{Median}(X)+r$ and $t=r/\sqrt{b}$ we get that
$\mathbb{P}\Big{(}X\geq\operatorname{Median}(X)+r\Big{)}\leq 2e^{-r^{2}/4d(\operatorname{Median}(X)+r)}\leq 2e^{-r/8d}+2e^{-r^{2}/8d\operatorname{Median}(X)}\,.$
@@
$\leq\int_{0}^{\infty}4\exp\Big{(}-\tfrac{t^{2}}{8d\operatorname{Median}(X)}\Big{)}+4\exp(-t)\mathrm{d}t\leq 10d\sqrt{\operatorname{Median}(X)}\,.$ (3.5)
@@
${\leq}\int_{n^{\delta+o(1)}}^{\infty}\exp\Big{(}n^{-1-\beta+2\alpha+o(1)}\big{(}n^{2\delta+o(1)}+tn^{\delta+o(1)}\big{)}\Big{)}\exp(-ct^{2})\mathrm{d}t$
@@
where the inequality follows from $\beta<\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$. Note that $12\leq|A_{\ell}|\leq d-$ 1 and $\delta<\tfrac{d-1}{d+1}$, we get that
@@
=\ = $-\tfrac{1}{d}\sum_{1\leq\ell\leq k}\big{(}|A_{\ell}|-d\big{)}\big{(}(1-\tfrac{\delta}{d-1})|A_{\ell}|-(1+\delta)\big{)}\geq 0\,,$
@@
We first prove Item (1). Define $\mu^{\prime}=\mu\big{(}\cdot\mid|\mathcal{H}|\leq(1+o(1))s\tbinom{n}{d}\big{)}$. Conditioned on $\mathcal{H}\sim\mu^{\prime}$, we define $\mathbb{P}^{\prime}$ by letting $\mathbb{P}^{\prime}(\cdot\mid\mathcal{H})\triangleq\mathbb{P}(\cdot\mid\mathcal{H})$. It is straightforward to check that $\operatorname{TV}(\mathbb{P}^{\prime},\mathbb{P})=o(1)$. Thus, by data processing inequality, it suffices to show that when $p=o(1),q=0$, for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
(|^$\mathbb{P}^{\prime}\Big{(}|\widehat{\mathcal{H}}(A)\cap\mathcal{H}|\geq\delta\cdot s\tbinom{n}{d}\Big{)}=o(1)\mbox{ for any constant }\delta>0\,.$
@@
For our purpose, for any $K=(1+o(1))s\tbinom{n}{d}$ we let $\mathcal{W}=\{\mathcal{H}\subset\tbinom{[n]}{d}:|\mathcal{H}|=K\}$. Noting that $\mathcal{H}\longrightarrow A\longrightarrow\widehat{\mathcal{H}}$ is a Markov chain and one can define a distance metric over $\mathcal{W}\times\mathcal{W}$ by taking $d(\mathcal{H},\mathcal{H}^{\prime})=\tfrac{|\mathcal{H}\triangle\mathcal{H}^{\prime}|}{2s\binom{n}{d}}$. For any constant $10\leq\epsilon<$ 1, we then have
@@
where $M=|\mathcal{W}|$ and $M_{\epsilon}=|\{\mathcal{H}^{\prime}\in\mathcal{W}:|\mathcal{H}^{\prime}\cap\mathcal{H}|\geq(1-\epsilon)s\binom{n}{d}\}|$. Clearly we have (recall that $K=n^{1+\delta+o(1)}$)
@@
Recall that in (2.6) we have shown that $\mu(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mu,\mu^{\prime})=o(1)$. Fix any $\mathcal{H}$ such that $|\mathcal{H}|=(1+o(1))s\tbinom{n}{d}$ and $|E(\mathcal{P}(\mathcal{H}))|=(1-o(1))s\tbinom{n}{d}\tbinom{d}{2}$, it is easy to check that with probability $1-o(1)$ we have
@@
To prove Item (2), without loss of generality, we assume $p=1$. A crucial observation (as shown in [BGPY25, Definition 1.6 and Lemma 2.5]) is that we only need to prove the following result: suppose we sample $(\mathcal{H},A)$ according to $\mathbb{P}^{\prime}$ and we sample $\mathcal{H}^{\prime}$ according to the posterior distribution of $\mathcal{H}$ given $A$. Then for all fixed $\iota=\Omega(1)$ we have

### v2

@@ Abstract
For a $d$-uniform random hypergraph on $n$ vertices in which hyperedges are included i.i.d. so that the average degree in the hypergraph is $n^{\delta+o(1)}$, the projection of such a hypergraph is a graph on the same $n$ vertices where an edge connects two vertices if and only if they belong to a same hyperedge. In this work, we study the inference problem where the observation is a noisy version of the graph projection where each edge in the projection is kept with probability $p=n^{-1+\alpha+o(1)}$ and each edge not in the projection is added with probability $q=n^{-1+\beta+o(1)}$. For all constant $d$, we establish sharp thresholds for both detection (distinguishing the noisy projection from an Erdős-Rényi random graph with edge density $q$) and reconstruction (estimating the original hypergraph). Notably, our results reveal a detection-reconstruction gap phenomenon in this problem. Our work also answers a problem raised in [BGPY25+].
@@ 1 Introduction
Community detection aims to partition a network’s vertices into clusters based on observed noisy associations between them. The most widely studied formalization assumes pairwise interactions: in friendship graphs, social networks, or co-purchase data, edges appear more frequently between vertices of the same latent community [New06]. However, many real-world systems are fundamentally characterized by higher-order interactions. In scientific co-authorship networks, for instance, relationships are intrinsically multi-way, involving groups of researchers rather than isolated pairs [New04]. Such data are naturally modeled by hypergraphs, a perspective that has led to the hypergraph stochastic block model (HSBM), a framework receiving considerable recent attention [GD14, ACKZ15, SZ22, GJ23, DW23, GP23, GP24].
This work focuses on a particular subproblem within this broad area: recovering multi-vertex association data (i.e., a hypergraph) from the pairwise association data (i.e., a graph). A motivating example is a large team collaborating on multiple projects, where the ground truth multi-way association is defined by project membership. However, due to practical limitations, we may only observe email exchanges between pairs of coworkers [KY04], which can be simplified to a graph where an edge connects any two individuals who communicated. Typically, such data is represented by constructing a graph from pairwise communications, effectively projecting the richer hypergraph onto a pairwise structure. Since such a projection generally loses information, a key question arises: to what extent can we reconstruct the original hypergraph from its (potentially noisy) graph projection, under the assumption that vertices within the same hyperedge exhibit higher edge density.
The problem of reconstruction from the projected graph has been extensively studied in recent years, both theoretically and empirically. For instance, [WK24] assumes access to a sample of a hypergraph from the underlying distribution and proposes a scoring method to select hyperedges based on their similarity to the sampled hypergraph. Another Bayesian algorithmic approach (i.e., to sample from the posterior distribution given the projected graph) was proposed in [YPP21, LYA23]. In addition, in [CFLL24] a large foundational model is used to recover a weighted hypergraph from a sample of its hyperedges, where each hyperedge is assigned a probability proportional to its weight. In a series of recent progress [BGP24, BGPY25+], the sharp threshold for reconstructing the hypergraph based on its projected graph in the noiseless setting was determined. They further raised the question of whether an analogous reconstruction threshold can be established in the noisy setting. Determining the reconstruction threshold in the noisy setting presents fundamental challenges absent in the noiseless case. In the noiseless model, the projected graph is a deterministic function of the latent hypergraph since every hyperedge manifests as a clique. The problem thus reduces to combinatorially identifying these cliques. In contrast, the noisy setting introduces two layers of uncertainty: hyperedges may fail to generate all their constituent pairwise edges (false negatives), while random background edges can connect vertices not sharing a hyperedge (false positives). This stochasticity destroys the clean clique structure, turning the problem into one of distinguishing subtle statistical deviations in edge density from a random baseline. Consequently, new techniques are required to prove sharp thresholds in the presence of this two-sided noise, making the problem significantly more intricate.
In this work, we not only settle the value of the reconstruction threshold for all $d$, but also establish the value of the (generally higher) threshold for detection, in which the algorithm only wants to detect the existence of group structure. In addition to finding the value of the threshold, we also establish a so-called “detection-reconstruction gap” for our problem, a phenomenon where the reconstruction task is strictly harder than detection. This gap, observed in other fundamental statistical problems [DJ04, MST19], is characterized by a distinct separation between the parameter regimes where detection is possible and where partial reconstruction becomes feasible.
2 Model definition and main results
We pose the model as follows: for any constant $d\in\mathbb{N}$, we denote by $\operatorname{U}=\operatorname{U}_{n,d}$ the collection containing all subsets of $n]=\{1,\ldots,n\}$ with cardinality $d$.
@@ 2 Model definition and main results
Define $\mathcal{H}$ to be the random $d$-uniform hypergraph on $[n]$ generated by including each $\Psi\in\operatorname{U}_{n,d}$ in $\mathcal{H}$ with probability $s=n^{-(d-1)+\delta+o(1)}$ for some constant $0\leq\delta<1$ [^1] For any two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write equivalently $a_{n}=O(b_{n})$, $b_{n}=\Omega(a_{n})$, $a_{n}\lesssim b_{n}$ and $b_{n}\gtrsim a_{n}$ if there exists a positive absolute constant $c$ such that $a_{n}/b_{n}\leq c$ holds for all $n$. We write $a_{n}=o(b_{n})$, $b_{n}=\omega(a_{n})$, $a_{n}\ll b_{n}$, and $b_{n}\gg a_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. We write $a_{n}=\Theta(b_{n})$ if both $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$ hold. Without further specification, all asymptotics are taken with respect to $n\to\infty$.. We denote the law of $\mathcal{H}$ by $\mu=\mu_{n,d,\delta}$. For simplicity, we will also use $\mathcal{H}$ to denote the set of hyperedges and $|\mathcal{H}|$ the cardinality of $\mathcal{H}$. In addition, for any $d$-uniform hypergraph $\mathcal{H}$, define $\mathcal{P}(\mathcal{H})$ to be the projection graph on $[n]$ where for all unordered pairs $(i,j)$ with $i\neq j$
@@
We require the condition $q\leq p$, reflecting the natural expectation that edges in $\mathcal{P}(\mathcal{H})$ tend to appear in the graph $A$. Two natural inference problems regarding this model are as follows: (1) the detection problem, i.e., testing $\mathbb{P}$ against $\mathbb{Q}=\mathbb{Q}_{n,q}$, where $\mathbb{Q}$ is the law of an Erdős-Rényi graph on $[n]$ with edge density $q=n^{-1+\beta+o(1)}$; (2) the reconstruction problem, i.e., estimating the original hypergraph $\mathcal{H}$ according to the observation $A(\mathcal{H})$.
In a series of recent progress [BGP24, BGPY25+], the sharp threshold for reconstructing the hypergraph $\mathcal{H}$ based on $A(\mathcal{H})$ in the noiseless setting $p=1,q=0$ was determined. In particular, it was shown in [BGPY25+] that when $\delta>\tfrac{d-1}{d+1}$ it is information theoretically impossible to reconstruct a positive fraction of $\mathcal{H}$ and when $\delta<\tfrac{d-1}{d+1}$ there exists an efficient algorithm that reconstructs a 1o(1)1-o(1) fraction of $\mathcal{H}$. They further raised the question of whether an analogous reconstruction threshold can be established in the noisy setting. Our main result can be informally summarized as follows:
@@
    (Almost-exact reconstruction) Suppose either $p=1-\Omega(1)$, $\delta>\tfrac{d-1}{d+1}$ or $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then it is informationally impossible to reconstruct a 1o(1)1-o(1) fraction of $\mathcal{H}$. Conversely, if $p=1-o(1),\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$, then there exists an efficient algorithm that successfully reconstructs a 1o(1)1-o(1) fraction of $\mathcal{H}$.
Notably, our result reveals a detection-reconstruction gap phenomenon in this problem where when $p=\Theta(1)$ (which implies that $\alpha=1$), detection is possible for all $\beta,\delta\in(0,1)$ but (partial) reconstruction is possible if and only if $\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. See Figure 1 for a summary of the regions where detection/reconstruction is possible. The reader may wonder whether the detection-recovery gap could be eliminated by choosing an alternative null distribution whose edge density matched that of the planted model $\mathbb{P}$. We show in Appendix A that this adjustment closes the gap partially but not entirely. In particular, detection remains possible at least whenever $\delta,\beta<\tfrac{d-1}{d+1}$, while for $\delta<\tfrac{d-1}{d+1}$ we always have $\tfrac{d-1}{d+1}>\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$.
Figure 1: Summary of the detection-reconstruction gap when $p=\Theta(1)$. Red regime II: reconstruction of $\mathcal{H}$ is possible; Blue regime IIII: detection with a better null distribution $\widetilde{\mathbb{Q}}$ (that matches the average edge-density) is possible, but reconstruction of $\mathcal{H}$ is impossible; Yellow regime IIIIII: detection with the simple null distribution $\mathbb{Q}$ is possible, but detection with a better null distribution $\widetilde{\mathbb{Q}}$ is unknown.
@@ 2.2 The reconstruction threshold
Recall the definition of $s$ in 2.1. A test statistic $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ is said to achieve
@@
$\widehat{\mathcal{H}}:=\big\{\Psi\in\operatorname{U}:A_{i,j}=1\mbox{ for all }i,j\in\Psi\big\}\,.$ (2.4)
@@
Suppose $\delta<\tfrac{d-1}{d+1},\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$ and p=1o(1)p=1-o(1). We have
@@
The proof of Proposition 2.7 is incorporated in Section 5. Next we show the impossibility of weak reconstruction. Note that when $\delta>\tfrac{d-1}{d+1}$, using [BGPY25+, Theorem 1.1] it is informationally impossible to achieve weak reconstruction even when $p=1,q=0$. Thus, without loss of generality, we assume that $\delta\leq\tfrac{d-1}{d+1}$.
@@
    Suppose that $p=o(1)$. Then for any estimator $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
    Suppose $q=n^{-1+\beta+o(1)}$ where $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then for any estimator $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
Suppose that $p=1-\Omega(1)$. Then for any estimator $\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
@@
$|E(A(\mathcal{H}))|\overset{d}{=}\mathbf{B}\big(|E(\mathcal{P}(\mathcal{H}))|,p\big)+\mathbf{B}\big(\tbinom{n}{2}-|E(\mathcal{P}(\mathcal{H}))|,q\big)\,.$ (3.1)
@@
Let $M_{k}(\mathcal{H})$ be the set of $(i,j)\in\mathcal{P}(\mathcal{H})$ such that there exist at least $k$ hyperedges in $\mathcal{H}$ that contain $(i,j)$. In addition, define
@@
which yields (3.2). Now conditioning on $\mathcal{P}(\mathcal{H})$ satisfying (3.2), by (3.1) and a standard binomial tail estimate, we obtain that with probability 1o(1)1-o(1) under $\mathbb{P}$
@@
Based on Lemma 4.2, it suffices to bound the tail probability of $|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|$, as shown in the following lemma.
@@
The proof of this inequality can be found in, e.g., [AS16, Section 7.7]. We now proceed to the proof of Lemma 4.3.
@@
where we view $(\mathcal{H},\mathcal{H}^{\prime})$ as a vector in $\{0,1\}^{2\binom{n}{d}}$. It is straightforward to verify that $h$ is Lipschitz. Furthermore, we claim that the function $h$ defined above is $f$-certifiable, where $f(s)=2\tbinom{d}{2}s\leq d^{2}s$. In fact, if $h(\mathcal{H},\mathcal{H}^{\prime})\geq s$ for some $(\mathcal{H},\mathcal{H}^{\prime})\in\Omega$, then there exists $E=\{(i,j):1\leq i<j\leq n\}$ such that for all $e\in E$ there exist $\zeta_{e},\zeta_{e}^{\prime}\in\tbinom{[n]}{d}$ with $\mathcal{H}(\zeta_{e})=\mathcal{H}^{\prime}(\zeta^{\prime}_{e})=1$. Denote $I=\{\zeta_{e}:e\in E\}\cap\{\zeta^{\prime}_{e}:e\in E\}$. Then for all $(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\in\{0,1\}^{2\binom{n}{d}}$ that agrees with $(\mathcal{H},\mathcal{H}^{\prime})$ on $I$, we have $)e\in E(\mathcal{P}(\widetilde{\mathcal{H}}))\cap E(\mathcal{P}(\widetilde{\mathcal{H}}^{\prime}$)) and thus $h(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\geq s$. As a result, Theorem 4.5 yields that for any $b,t\geq 0$,
$\mathbb{P}\big(X\leq b-td\sqrt{b}\big)\cdot\mathbb{P}(X\geq b)\leq e^{-t^{2}/4}\,.$
@@
$\mathbb{P}\Big(X\leq\operatorname{Median}(X)-t\Big)\leq 2e^{-t^{2}/4d^{2}\operatorname{Median}(X)}\,.$
In addition, taking $b=\operatorname{Median}(X)+r$ and $t=r/d\sqrt{b}$ we get that
$\mathbb{P}\Big(X\geq\operatorname{Median}(X)+r\Big)\leq 2e^{-r^{2}/4d^{2}(\operatorname{Median}(X)+r)}\leq 2e^{-r/8d^{2}}+2e^{-r^{2}/8d^{2}\operatorname{Median}(X)}\,.$
@@
$\leq\int_{0}^{\infty}4\exp\Big(-\tfrac{t^{2}}{8d^{2}\operatorname{Median}(X)}\Big)+2\exp\big(-\tfrac{t}{8d^{2}}\big)\mathrm{d}t=4\sqrt{2\pi}d\sqrt{\operatorname{Median}(X)}+16d^{2}.$ (4.5)
@@
[…]
@@
where the inequality follows from $\beta<\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$. Note that $2\leq|A_{\ell}|\leq d-1$ and $\delta<\tfrac{d-1}{d+1}$, we get that
@@
$=\$$-\tfrac{1}{d}\sum_{1\leq\ell\leq k}\big(|A_{\ell}|-d\big)\big((1-\tfrac{\delta}{d-1})|A_{\ell}|-(1+\delta)\big)\geq 0\,$,
@@
We first prove Item (1). Note that if there is an estimator $\widehat{\mathcal{H}}$ such that $\mathbb{E}_{\mathbb{P}}[|\mathcal{H}\triangle\widehat{\mathcal{H}}|]=(1-\Omega(1))s\tbinom{n}{d}$. Then clearly we have
$\mathbb{P}\Big(|\widehat{\mathcal{H}}\cap\mathcal{H}|\geq\Omega(1)s\tbinom{n}{d},|\widehat{\mathcal{H}}|\leq(2-\Omega(1))s\tbinom{n}{d}\Big)\geq\Omega(1)\,.$
[…]
$\mathbb{P}^{\prime}\Big(|\widehat{\mathcal{H}}(A)\cap\mathcal{H}|\geq\delta\cdot s\tbinom{n}{d}\Big)=o(1)\mbox{ for any constant }\delta>0\,.$
@@
For our purpose, for any $K=(1+o(1))s\tbinom{n}{d}$ we let $\mathcal{W}=\{\mathcal{H}\subset\tbinom{[n]}{d}:|\mathcal{H}|=K\}$. In addition, since adding or deleting an arbitrary $o(K)$ elements of $\widehat{\mathcal{H}}$ does not affect the result, we can assume that $\widehat{\mathcal{H}}$ has cardinality $K$. Noting that $\mathcal{H}\longrightarrow A\longrightarrow\widehat{\mathcal{H}}$ is a Markov chain and one can define a distance metric over $\mathcal{W}\times\mathcal{W}$ by taking $d(\mathcal{H},\mathcal{H}^{\prime})=\tfrac{|\mathcal{H}\triangle\mathcal{H}^{\prime}|}{2s\binom{n}{d}}$. For any constant $0\leq\epsilon<1$, we then have
@@
[…]
@@
Recall that in (3.6) we have shown that $\mu(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mu,\mu^{\prime})=o(1)$. Fix any $\mathcal{H}$ such that $|\mathcal{H}|=(1+o(1))s\tbinom{n}{d}$ and $|E(\mathcal{P}(\mathcal{H}))|=(1-o(1))s\tbinom{n}{d}\tbinom{d}{2}$, it is easy to check that with probability 1o(1)1-o(1) we have
@@
To prove Item (2), without loss of generality, we assume $p=1$. A crucial observation (as shown in [BGPY25+, Definition 1.6 and Lemma 2.5]) is that we only need to prove the following result: suppose we sample $(\mathcal{H},A)$ according to $\mathbb{P}^{\prime}$ and we sample $\mathcal{H}^{\prime}$ according to the posterior distribution of $\mathcal{H}$ given $A$. Then for all fixed $\iota=\Omega(1)$ we have